OpenStudy (anonymous):

evaluate: cot(sec inverse x+sin inverse(1/X))

OpenStudy (anonymous):

\[\cot [\sec^{-1}(x) + \sin^{-1}(\frac{1}{x})]\] Is this what the question is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

prove it @Siddy

OpenStudy (anonymous):

can we use calculus?

OpenStudy (anonymous):

yeh @mukushla

OpenStudy (anonymous):

thts not true @Siddy

OpenStudy (anonymous):

cot(pi/2) = 0

OpenStudy (anonymous):

\[\cot(\sec^{-1}x + cosec^{-1}x ) \implies \cot(\frac{\pi}{2}) = 0\]

OpenStudy (anonymous):

let \[f(x)=arcsec \ x+ \arcsin \ \frac{1}{x} \\ f'(x)=\frac{1}{\left| x \right| \sqrt{x^2-1}}-\frac{1}{\left| x \right| \sqrt{x^2-1}}=0 \] so f(x) is constant

OpenStudy (anonymous):

\[\sin^{-1} \frac{1}{x} = cosec^{-1}x\]

OpenStudy (anonymous):

\[\sin^{-1}x + \cos^{-1} x = \frac{\pi}{2}\] \[\sec^{-1}x + cosec^{-1} x =\frac{\pi}{2}\] \[\tan^{-1}x + \cot^{-1} x = \frac{\pi}{2}\]

OpenStudy (anonymous):

Using these formula the term in the bracket can be changed to \(\sec^{-1}x + \csc^{-1} x = \frac{\pi}{2}\) and you will get your answer..

OpenStudy (anonymous):

will the answer be 0 ?

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

fine.thanks.

OpenStudy (anonymous):

Welcome..